OpenStudy (anonymous):

In 1987 a recording artist sold 14,000 albums. In 1998 that same artist sold 4,683 albums. Find the rate of change in record sales from 1987 to 1998.

OpenStudy (anonymous):

rate = (new-old)/(end-start) = (4683-14000)/(1998-1987) roughly a thousand per year decline....calculate exact #s

OpenStudy (anonymous):

so 847 albums per year?